# Elasticity without a reference state: continuum mechanics of active tension nets

### v2

@@ Discussion
While we have introduced our theory in the context of epithelial tissue mechanics, we believe it applies more broadly. Non-epithelial tissues, like those composed of protrusive, mesenchymal cells, do not form tension nets in the strict ATN sense, but their architecture and shape are also controlled by balancing active contractile stresses against bulk elasticity. Beyond the biological setting, 2D fluid or ferro-fluid foams [50, 13] correspond to the limiting case where all $\tau_{ij}=\tau=\mathrm{const}$. Tension dynamics could be introduced, for instance, through chemical reactions within foam bubbles. As explained in the companion paper [15], there is also a close connection to granular materials and truss networks, which can be understood as the mechanical Legendre dual of tension nets. In granular materials, one typically studies the ensemble of compatible (force-balanced) stress configurations for a given physical arrangement of particles. This is, in a sense, the dual of the problem considered in the present study. In future work, the Legendre duality might, therefore, provide a starting point for building a continuum theory for granular matter. Indeed, previous phenomenological approaches postulate that granular materials can also be described by “emergent elasticity” [51].
To conclude, we believe that our theory provides a general framework for the emergent elasticity and plasticity of matter whose microscopic mechanics is not governed by conventional constitutive laws but by independently controlled active forces. It quantitatively explains how these local active forces generate large-scale shape.
@@ Acknowledgements.
Appendix A Change of curvature radius by a conformal map
@@
Here, we compute the change of a line’s curvature radius $R$ by a conformal map $z\mapsto f(z)$ [40]. Consider a line $z(t)$ with normal $\mathbf{n}$}. The rotation of the local normal is given by the vorticity $\omega=\arg\partial_{z}f=\partial_{z}f/|\partial_{z}f|$. Moving a step $dt$ along the transformed curve $f(z(t))$, the tangent $i\mathbf{n}$ rotates by an angle $d\phi=\omega(f(z(t+dt))-\omega(f(z(t))=(i{\mathbf{n}})\cdot 2\bar{\partial}_{z}\omega dt$. Due to the scaling factor $|\partial_{z}f|$, the arc length is $|\partial_{z}f|dt$. Putting this together, the curvature radius reads:
$\frac{1}{R}$ $=\frac{d\phi}{d(\lambda t)}=\frac{1}{|\partial_{z}f|}(i{\mathbf{n}})\cdot 2\bar{\partial}_{z}\arg\partial_{z}f$
$={\mathbf{n}}\cdot 2\bar{\partial}_{z}|\partial_{z}f|$ (59)
Appendix B The isogonal-conformal decomposition
@@
Note that the curl of the LHS is $\epsilon_{bc}\partial_{\gamma}\partial_{\alpha}\partial_{\beta}\theta^{(2)}=0$. For Eq. (60) to be solvable, the curl of the RHS must also vanish:
@@
Appendix C Boundary conditions
@@
To fulfill $\partial_{n}\theta=u_{n}$ we continue $\theta$ into the bulk with slope $u_{n}$, i.e.
@@
As an example, consider the unit disk in complex polar coordinates $z=re^{i\phi}$. Conformal maps are holomorphic functions $f$ and, hence, analytic, $f(z)=\sum_{n\geq 0}\tfrac{f^{(n)}}{n!}z^{n}$. Importantly, the sum is restricted to $n\geq 0$: hence, the boundary value $f(e^{i\phi})$ has only non-negative Fourier coefficients. A boundary displacement with negative Fourier coefficients cannot be extended to a conformal deformation of the disk. For example, pure shear $\mathbf{u}|_{\partial\Omega}\propto e^{-i\phi}=x\hat{\mathbf{x}}-y\hat{\mathbf{y}}$ is incompatible with a conformal deformation.
Boundary conditions for the Beltrami equation
In the context of the Beltrami equation Eq. (10), the BCs fix the conformal contribution to $z(\xi,\bar{\xi})$. Therefore, the Beltrami equation also admits a single BC only, e.g., the natural BC for $|\partial_{\xi}z|$, Eq. (14). Indeed, the example Appendix G.1 makes it clear that it is impossible to make the tension metric isotropic while keeping the boundaries clamped. The Beltrami equation must be contrasted with the elasticity problem for the physical configuration $w$, which admits two boundary conditions, as we showed above.
Appendix D Adiabatic approximation
In the main text, we treat tissue dynamics within the adiabatic approximation. Consider relaxation dynamics of the cell positions $\mathbf{r}$
where $\gamma$ is a friction coefficient. The stress $\sigma$ depends both on $\mathbf{r}(t)$ and $\mathbf{g}(t)$ by Eq. (4), and is thus time-dependent. The timescale of relaxation to equilibrium is given by $t_{\gamma}=L\gamma/p_{0}$, where $L$ is the system length scale. In the limit where $t_{\gamma}$ is much faster than any other timescale (notably, that of tension dynamics $\partial_{t}\mathbf{g}$), Eq. (66) reduces to force balance:
$\partial_{\alpha}\sigma_{\alpha\beta}(t)-\partial_{\beta}p(t)=0$ (67)
The cells, therefore, adopt the instantaneous equilibrium configuration $\mathbf{w}({\bm{\xi}},t)$, whose time-dependence results from the dynamics of the “parameter” $\mathbf{g}(t)$.
Appendix E Integrating the Beltrami flow via Green’s functions
The Beltrami flow Eq. (40) describes the flow $v(z,\bar{z})$ of the stress-free reference state $z$ due to the dynamics of the tension metric. Note that for any solution $v(z,\bar{z})$ of Eq. (40), $v(z,\bar{z})+f(z)$ for a conformal $f(z)$ is also a solution. The BCs fix this conformal component. In the main text, we used natural BCs, $\lambda_{g}|_{\partial\Omega}=1$. This minimizes the deformation of the embedding $z(\zeta,\bar{\zeta})$ and the pressure gradients in the reference state, providing a convenient starting point for linearization (Eq. (.2.4)). The shape of the natural domain boundary $\partial\Omega$ is, however, time-dependent.
@@ Appendix E Integrating the Beltrami flow via Green’s functions
Eq. (68) also allows the explicit computation of the topological flow due to local cell rearrangement. In summary, the dynamics of the reference state becomes a QC flow $v_{(\hat{z})}$ on the disk $D$, and is analytically tractable. The price to be paid is that the deformation from the reference to the physical state is now, in general, very large.
@@
Denoting the conformal factor by $\lambda_{\hat{z}|z}=|\partial_{\hat{z}}z|$, the natural BCs require $\lambda_{\hat{z}|z}^{2}|_{\partial_{D}}=\lambda_{g}^{2}(\hat{z},\bar{\hat{z}})|_{\partial_{D}}$ on the disk boundary. We now note that $\Delta\log\lambda_{\hat{z}|z}=0$, since $\log\lambda_{\hat{z}|z}$ is the real part of the analytic function $\log(\partial_{\hat{z}}z)$. Therefore, we can obtain $\log\lambda_{\hat{z}|z}$ on the entire disk by convolution with the Poisson kernel $P(\hat{z},\hat{z}^{\prime})$ as
@@
In particular, the normal derivative on the boundary is $\partial_{\mathbf{n}}\log\lambda_{\hat{z}|z}|_{\partial D}=(2\pi)^{-1}\int_{\partial D}d\hat{z}^{\prime}\log\lambda_{\hat{z}|z}(\hat{z}^{\prime})\operatorname{Re}\left[\frac{\hat{z}\hat{z}^{\prime}}{(\hat{z}-\hat{z}^{\prime})^{2}}\right]$. Together, $\lambda_{\hat{z}|z}|_{\partial D}$ and $\partial_{\mathbf{n}}\lambda_{\hat{z}|z}|_{\partial D}$ fully describe the shape of the boundary: $\lambda_{\hat{z}|z}$ is the local arc length, and $\kappa=\lambda_{\hat{z}|z}^{-1}+\partial_{\mathbf{n}}\lambda_{\hat{z}|z}^{-1}$ is the curvature (Appendix A). The orientation of the tangent to the boundary at arc length $s$ is given by $\phi(s)=\int_{0}^{s}\kappa(s^{\prime})ds^{\prime}$. One can then integrate to obtain the boundary curve $z(\hat{z})|_{\partial D}$. Finally, the value in the interior of the disk is determined by Cauchy’s integral:
@@
Appendix F Kinematics of Beltrami flow
F.1 Convective derivative of $\mu_{\zeta|z}$
Consider the Beltrami flow $\zeta^{\prime}=\zeta+{\dot{\zeta}}dt$ under dynamics of the adjacency metric; $\bar{\partial}_{\zeta}{\dot{\zeta}}=\partial_{t}\mu_{a}$. We would like to calculate how $\mu_{\zeta|z}$ changes under this flow while holding the tension metric $\mathbf{g}$ fixed. To find the new Beltrami coefficient $\mu_{\zeta^{\prime}|z}$, we need to compose the maps $\zeta^{\prime}\mapsto\zeta$ and $\zeta\mapsto z$. Drawing the map compositions as a diagram

### v3

@@ Discussion
While we have introduced our theory in the context of epithelial tissue mechanics, we believe it applies more broadly. Non-epithelial tissues, like those composed of protrusive, mesenchymal cells, do not form tension nets in the strict ATN sense, but their architecture and shape are also controlled by balancing active contractile stresses against bulk elasticity. Beyond the biological setting, 2d fluid or ferro-fluid foams [61, 12] correspond to the limiting case where all $\tau_{ij}=\tau=\mathrm{const}$. Tension dynamics could be introduced, for instance, through chemical reactions within foam bubbles. As explained in the companion paper [18], there is also a close connection to granular materials and truss networks, which can be understood as the mechanical Legendre dual of tension nets. In granular materials, one typically studies the ensemble of compatible (force-balanced) stress configurations for a given physical arrangement of particles. This is, in a sense, the dual of the problem considered in the present study. In future work, this Legendre duality might provide a starting point for continuum theories for granular matter. Indeed, previous phenomenological approaches postulate that granular materials can also be described by “emergent elasticity” [62]. Lastly, one may ask if the tension-metric-based approach can be generalized from tissue sheets to three-dimensional “volumetric” tissues. In 3d, Riemannian geometry becomes considerably richer [24].
To conclude, we believe that our theory provides a general framework for the emergent elasticity and plasticity of active solids, matter whose microscopic mechanics is not governed by conventional constitutive laws but by independently controlled active forces. It quantitatively explains how these local active forces generate large-scale shape.
@@ Acknowledgements.
Appendix A Discrete-continuum correspondence for a “rectangular foam”
Figure 5: “Rectangular foam” with anisotropic surface tensions $\tau_{1}$ on horizontal edges and $\tau_{2}$ on vertical edges.
Here, we discuss a minimal “cartoon” example that connects the continuum theory to the discrete active tension network model. It generalizes the “square foam” introduced in [64] and is composed of identical rectangular cells with internal pressure $p$ and possibly anisotropic interfacial tensions $\tau_{1},\tau_{2}$ on horizontal and vertical edges (see Fig. 5.) Batchelor’s formula [65] (treating each edge as a force dipole) gives the macroscopic tensile stress
$\bm{\sigma}=\frac{1}{a}\sum_{i}\ell_{i}\tau_{i}\,\hat{\mathbf{e}}_{i}\otimes\hat{\mathbf{e}}_{i}=\frac{1}{\ell_{1}\ell_{2}}\begin{pmatrix}\tau_{1}\ell_{1}&0\\ 0&\tau_{2}\ell_{2}\end{pmatrix}$ (64)
where $\ell_{i}$ are the edge lengths. In an unloaded state (no external forces), force balance implies $0=dE=\tau_{1}d\ell_{1}+\tau_{2}d\ell_{2}-pd(\ell_{1}\ell_{2})$ and therefore $\ell_{1,2}^{0}=\tau_{2,1}/p$. Hence $\sigma_{\alpha\beta}=p\delta_{\alpha\beta}$ and the total stress $\sigma^{\mathrm{tot}}=\sigma_{\alpha\beta}-p\delta_{\alpha\beta}$ vanishes identically, as expected in an unloaded uniform state. It is worth emphasizing that the configuration $\ell^{0}_{1,2}$ is an effective stress-free reference state: the tensions $\tau_{1},\tau_{2}$ are length-independent. Deforming the system away from this reference state leads to a non-zero total stress
$\bm{\sigma}^{\mathrm{tot}}\approx-p\begin{pmatrix}\frac{\ell_{2}}{\ell^{0}_{2}}-1&0\\ 0&\frac{\ell_{1}}{\ell^{0}_{1}}-1\end{pmatrix}$ (65)
Recognizing the elements of the strain tensor, we read off the effective shear and bulk moduli $\mu_{\mathrm{eff}}=p/2$ and $B_{\mathrm{eff}}=-p/2$. Thus, the elastic response is evidently isotropic, despite the underlying microscopic anisotropy. Moreover, the negative bulk modulus indicates mechanical instability. This instability is a consequence of simultaneously prescribing edge tensions and pressure. It implies that pressure cannot be prescribed, but must derive from an equation of state $p=P(n)$, where $n=1/(\ell_{1}\ell_{2})$ is the cell density. To find the unloaded state, one needs to solve self-consistently for $p_{0}=P(n_{0})=P(p_{0}^{2}/\tau_{1}\tau_{2})$. The linearization of the equation of state around $n_{0}$ then yields an additional contribution to the effective bulk modulus $B_{\mathrm{eff}}=-\frac{p_{0}}{2}+P^{\prime}(n_{0})$, restoring stability for sufficiently large $P(n_{0})>p_{0}/2$.
We conclude our discussion by illustrating the tension metric and its isothermal embedding. Consider Lagrangian coordinates $\bm{\xi}$ enumerating rows and columns of cells. In these coordinates, we can encode the edge tensions in the tension metric
$\mathbf{g}(\bm{\xi})=\begin{pmatrix}\tau_{2}^{2}&0\\ 0&\tau_{1}^{2}\end{pmatrix}$. (66)
Note that $\tau_{2}$ appears in the $(1,1)$ entry since the tension metric encodes tension on edges perpendicular to a displacement vector $d\bm{\xi}$. In the physical configuration, specified by $\ell_{1},\ell_{2}$, the metric reads
$\mathbf{g}(\mathbf{r})=\begin{pmatrix}\tau_{2}^{2}/\ell_{1}^{2}&0\\ 0&\tau_{1}^{2}/\ell_{2}^{2}\end{pmatrix}$. (67)
This ensures that $\tau_{i}^{2}=g_{\alpha\beta}\,dr_{i,\alpha}dr_{i,\beta},\;i=1,2$, where $d\mathbf{r}_{i}$ is the physical vector connecting two adjacent cells. Crucially, comparing Eqs. (64) and (67) recovers the stress-metric relation Eq. (4):
$\bm{\sigma}^{2}(\mathbf{r})=\begin{pmatrix}\tau_{1}^{2}/\ell_{2}^{2}&0\\ 0&\tau_{2}^{2}/\ell_{1}^{2}\end{pmatrix}=\bm{\epsilon}^{T}\!\cdot\mathbf{g}(\mathbf{r})\cdot\bm{\epsilon}$ (68)
To find the isothermal embedding where $\mathbf{g}\propto\mathbb{I}$, one thus scales the spatial directions by the respective tensions: $p_{0}\mathbf{r}_{1}=\tau_{2}\bm{\xi}_{1}$, $p_{0}\mathbf{r}_{2}=\tau_{1}\bm{\xi}_{2}$. This recovers the stress-free reference calculated above.
This toy example already captures key elements of actual tension networks and foams, which are locally hexagonal instead of rectangular, and whose tension configuration is not spatially uniform. The calculation of the effective stress of these more complex networks is a key part of the companion paper [18].
Appendix B Examples for the static problem
We provide two concrete examples to illustrate the continuum equations derived in the main text. Since the static problem is mathematically equivalent to linear elasticity, these examples serve to clarify the definitions of the tension- and adjacency metrics and the logic of the mapping from the tension manifold to the physical state.
B.1 Statics: flat tension metric
As a first example, we revisit the rectangular foam from App. A now in the continuum, as illustrated in Fig. 6. Let the adjacency metric be the Euclidean metric, corresponding to uniformly distributed isotropic cells. The $a$-isothermal $\zeta$ coordinates are therefore simply Cartesian coordinates in the unit square $0\leq\zeta_{1},\zeta_{2}\leq 1$. In this square shape, the cell density is uniform, and cell shapes are isotropic. Next, we need to specify the active tension configuration, i.e., the tension metric $g$. In this example, we consider an anisotropic metric, $g_{\zeta\zeta}=2,\quad g_{\zeta\bar{\zeta}}=s$ where the real parameter $0<s\ll 1$ determines the strength of anisotropy. In Cartesian notation, $\mathbf{g}(\zeta_{1},\zeta_{2})=\mathrm{diag}[1+s,1-s]$ corresponding to $\tau_{1}^{2}=1-s$, $\tau_{2}^{2}=1+s$ in the notation of App. A. Microscopically, this tension metric means that vertical cell-cell interfaces are under higher tension than horizontal ones, since the tension metric encodes the tension on interfaces transverse to a line element $d\mathbf{r}$ (cf. Fig. 2). The Beltrami coefficient reads $\mu_{g}=s/(2+\sqrt{(1+s)(1-s)})\approx s/4$, where the approximation is for $s\ll 1$. To remove this anisotropy, the $\mathbf{g}$-isothermal coordinates need to create a corresponding shear, $z(\zeta,\bar{\zeta})=\zeta+\mu_{g}\bar{\zeta}$. This fulfills the Beltrami Eq. (10). In Cartesian notation, $\mathbf{z}=\mathrm{diag}[1+\mu_{g},1-\mu_{g}]\cdot\bm{\zeta}$.
Figure 6: Elementary example for tension- and adjacency metrics.
Since $\det\mathbf{g}\approx 1$, tension and cell densities are (approximately) equal in this example. Further, since $\mathbf{g}$ is constant, it has no curvature, $K_{g}=0$: the $z$-configuration has constant cell density and constant pressure $p_{0}$. It is therefore mechanically balanced in the absence of external forces. In other words, $u_{\alpha}=0$ solves the linear elasticity force-balance equation Eq. (32). This example illustrates how tension anisotropy controls tissue shape.
Let us now consider applying a constant boundary traction force $b_{1}=0,b_{2}=\pm b$ to the upper (++) and lower (-) edges of the rectangular tissue patch, (the left and right sides remain free). This results in a potential displacement $u_{\alpha}=f_{\alpha}+\partial_{\alpha}\theta$ from the tension reference $z$ to the physical configuration $w=z+u(z,\bar{z})$. By Eq. (35), the potential $\theta$ must be biharmonic, $\Delta^{2}\theta=0$. The traction forces $\sigma\cdot\mathbf{n}_{\pm}=[0,\pm b]$ at the top and bottom edges require $\sigma_{22}|_{\pm}=p_{0}(\partial_{1}^{2}\theta)|_{\pm}=b$. We can hence take $\theta=b\operatorname{Re}[z]^{2}/(2p_{0})$, which implies a uniform stress $\bm{\sigma}=\mathrm{diag}\,(0,b)$ in the bulk and, thus, also fulfills the free BCs on the left and right edge. This result is expected by simple physical reasoning. To ensure that the displacement $u_{\alpha}$ is incompressible, we add a conformal contribution f(z)=bzf(z)=-bz (in this case, simply a uniform scaling).
B.2 Statics: non-flat tension metric
Next, we study an example where the tension metric has curvature. We consider a circular tissue patch. The adjacency metric is the Euclidean metric in the unit disk. The $\mathbf{a}$-isothermal $\zeta$-coordinates are simply Cartesian coordinates. In $\zeta$-coordinates, the cell density is uniform. We now define the tension metric by:
We now need to find the transformation $z(\zeta,\bar{\zeta})$ that makes this tension metric isotropic. Because of rotation symmetry, we expect a radial dilation. The $g$-isothermal coordinates are given by the stereographic projection [66] of the hemisphere, which is a conformal map:
$z(\zeta,\bar{\zeta})=\frac{\zeta}{1+h(\zeta,\bar{\zeta})}$ (70)
Due to curvature, area distortion (and thus cell density gradients) in the tension reference state $z$ are unavoidable. For an incompressible tissue, these must be compensated by an isogonal displacement $w-z=u=f(z)+2\bar{\partial}_{z}\theta$; Eq. (32). Incompressibility means $(\det\mathbf{D}w)^{-1}\sqrt{\det g(z,\bar{z})}=1$ where $(Dw)_{\alpha\beta}=\partial_{\alpha}w_{\beta}$ is the Jacobian. Expanding $\det\mathbf{D}w=\det[\delta_{\alpha\beta}+\partial_{\alpha}u_{\beta}]$ with $w_{\alpha}=f_{\alpha}+\partial_{\alpha}\theta$ to linear order:
$\Delta\theta+\partial_{\alpha}f_{\alpha}=\frac{2}{z\bar{z}-1}$ (71)
$\Delta\theta(r)+C=\frac{2}{1+r^{2}}$ (72)
with BCs $\theta(0)=0$, $\partial_{r}\theta(r=0)=0$, and $\partial_{r}\theta(r=0)=0$. The last condition follows from $\sigma\cdot\mathbf{n}=0$ (vanishing traction forces at the boundary). Via the strain $C=\partial_{\alpha}f_{\alpha}$ of the conformal map, we fulfill the last BC. Equqtion (72) is solved by $\theta(r)=\tfrac{1}{4}[-r^{2}+2\log(1+r^{2})]$ with $C=1$.
Within the disk, $\sigma_{\alpha\beta}^{\mathrm{tot}}=-p_{0}\epsilon_{\alpha\gamma}\epsilon_{\beta\delta}\partial_{\gamma}\partial_{\delta}\theta\neq 0$; Eq. (34). These internal residual stresses, shown in Fig. 7, result from curvature of the tension manifold. Importantly, such stresses will also be present for finite cell compressibility. In the limit of fully compressible cells, no force-balanced configuration exists because sustaining the pressure field implied by the tension manifold curvature requires a finite cell compressibility.
Appendix C Change of curvature radius by a conformal map
@@
Here, we compute the change of a line’s curvature radius $R$ by a conformal map $z\mapsto f(z)$ [66]. Consider a line $z(t)$ with normal $\mathbf{n}$}. The rotation of the local normal is given by the vorticity $\omega=\arg\partial_{z}f$. Moving a step $dt$ along the transformed curve $f(z(t))$, the tangent $i\mathbf{n}$ rotates by an angle $d\phi=\omega(f(z(t+dt)))-\omega(f(z(t)))=(i{\mathbf{n}})\cdot 2\bar{\partial}_{z}\omega dt$. Due to the scaling factor $|\partial_{z}f|$, the arc length is $|\partial_{z}f|dt$. Altogether, the curvature radius reads:
$\frac{1}{R}$ $=\frac{d\phi}{d(\lambda t)}=\frac{i{\mathbf{n}}}{|\partial_{z}f|}\cdot 2\bar{\partial}_{z}\arg\partial_{z}f={\mathbf{n}}\cdot 2\bar{\partial}_{z}|\partial_{z}f|$ (73)
Appendix D The isogonal-conformal decomposition
@@
Since the curl of the LHS is $\epsilon_{\beta\gamma}\partial_{\gamma}\partial_{\alpha}\partial_{\beta}\theta^{(2)}=0$, the curl of the RHS must also vanish:
@@
Appendix E Boundary conditions
@@
To fulfill $\partial_{n}\theta=u_{n}$, we continue $\theta$ into the bulk with slope $u_{n}$:
@@
For example, consider the unit disk in complex polar coordinates $z=re^{i\phi}$. Conformal maps $f$ are analytic, $f(z)=\sum_{n\geq 0}\tfrac{f^{(n)}}{n!}z^{n}$. Importantly, the sum is restricted to $n\geq 0$: hence, the boundary value $f(e^{i\phi})$ has only non-negative Fourier coefficients. A boundary displacement with negative Fourier coefficients cannot be extended to a conformal deformation of the disk. For example, pure shear $\mathbf{u}|_{\partial\Omega}\propto e^{-i\phi}=x\hat{\mathbf{x}}-y\hat{\mathbf{y}}$ is incompatible with a conformal deformation.
Boundary conditions for the Beltrami equation.
In the context of the Beltrami equation Eq. (10), the BCs fix the conformal contribution to $z(\xi,\bar{\xi})$. Therefore, the Beltrami equation also admits a single BC only, e.g., the natural BC for $|\partial_{\xi}z|$, Eq. (14). Indeed, the example Appendix B.1 makes it clear that it is impossible to make the tension metric isotropic while keeping the boundaries clamped. The Beltrami equation must be contrasted with the elasticity problem for the physical configuration $w$, which admits two boundary conditions.
Appendix F Integrating the quasiconformal flow via Green’s functions
The quasiconformal flow Eq. (42) describes the flow $\dot{z}(z,\bar{z})$ of the stress-free reference state $z$ due to the dynamics of the tension metric. Note that for any solution $\dot{z}(z,\bar{z})$ of Eq. (42), $\dot{z}(z,\bar{z})+f(z)$ for a conformal $f(z)$ is also a solution. The BCs fix this conformal component. In the main text, we used natural BCs, $\lambda_{g}|_{\partial\Omega}=1$. This minimizes the deformation of the embedding $z(\zeta,\bar{\zeta})$ and the pressure gradients in the reference state, providing a convenient starting point for linearization; Eq. (.2.4). The shape of the natural domain boundary $\partial\Omega$ is, however, time-dependent.
@@ Appendix F Integrating the quasiconformal flow via Green’s functions
Eq. (80) also allows the explicit computation of the topological flow due to local cell rearrangement. In summary, the dynamics of the reference state becomes a QC flow $\dot{\hat{z}}$ on the disk $D$, and is analytically tractable. The price to be paid is that the deformation from the reference to the physical state is now, in general, very large.
@@
In particular, the normal derivative on the boundary is $\partial_{\mathbf{n}}\log\lambda_{\hat{z}|z}|_{\partial D}=(2\pi)^{-1}\int_{\partial D}d\hat{z}^{\prime}\log\lambda_{\hat{z}|z}(\hat{z}^{\prime})\operatorname{Re}\left[\frac{\hat{z}\hat{z}^{\prime}}{(\hat{z}-\hat{z}^{\prime})^{2}}\right]$. Together, $\lambda_{\hat{z}|z}|_{\partial D}$ and $\partial_{\mathbf{n}}\lambda_{\hat{z}|z}|_{\partial D}$ fully describe the shape of the boundary: $\lambda_{\hat{z}|z}$ is the local arc length, and $\kappa=\lambda_{\hat{z}|z}^{-1}+\partial_{\mathbf{n}}\lambda_{\hat{z}|z}^{-1}$ is the curvature (Appendix C). The orientation of the tangent to the boundary at arc length $s$ is given by $\phi(s)=\int_{0}^{s}\kappa(s^{\prime})ds^{\prime}$. One can then integrate to obtain the boundary curve $z(\hat{z})|_{\partial D}$. Finally, the value in the interior is determined by Cauchy’s integral:
@@
Appendix G Kinematics of quasiconformal flow
G.1 Convective derivative of $\mu_{\zeta|z}$
Consider the quasiconformal flow $\zeta^{\prime}=\zeta+{\dot{\zeta}}dt$ under dynamics of the adjacency metric; $\bar{\partial}_{\zeta}{\dot{\zeta}}=\partial_{t}\mu_{a}$. We calculate how $\mu_{\zeta|z}$ changes under this flow while holding the tension metric $\mathbf{g}$ fixed. To find the new Beltrami coefficient $\mu_{\zeta^{\prime}|z}$, we need to compose the maps $\zeta^{\prime}\mapsto\zeta$ and $\zeta\mapsto z$. Drawing the map compositions as a diagram